4. Solve the IVP y + 4y = 36t² + 34t, y(0) = 0, y0) = 0 b) 4y - y= 4 + 122, y(0) = 0, 7(0) = 0, y0) = 0
Answer #1

y+qy = 3642 +3 at y lolzo Y(0) So the characterstic Creation is gut qy=o 2 +40=o o D(0+4)=0 DEO, - complementy solution Y9р - 342 424 о (не) OPE ( 2 ) 1 ст за 26E1 11q Эр- І, 1 - D4 D 36 +1+34 +1 to takes ) ! (at 4.4 16 p = 3, E414 - 16+ – 65 1The complete solution al chifferential eguation y = yetup y = atzet + 3t²-24²_t Apply the boundary condition as to y=0 o=ataqy - ye at +12624 y(0) 0 7 (0)=0 y (o)=0 The characterstic @geahon is 40²-0=0 D(?p2-03 59 0-0, 40-70 D² = 19 7 D= I12 compYo = qt + 12 e 2t 9.D ² P yp- 4 + 12e2t tt 12024 D(40²-1) -D.(1-482) Y p = (1-2025 ( 44 +12224) yoon+ (1+202) (94 +1264) ypeThen the saletion of o the the differ differential care 9 - 9 + 9р 9= 0.5t -0.5t Citae tge a 2t²_102e21 apply the boundary cso= 0.25 G +0.25 - 4-208 0.25 (2 to 250g= a12 czteaz 162€ - ☺ C2 teg= 1696 Solac eg © e 262 = 1646 2 = 2 = 829) C2 = 624 from